OpenStudy (kamille):

Hello! Maybe someone has some time to help me? So, with which values of "a" equatation of "ax+2=x" solution is less than 3?

OpenStudy (tamtoan):

subtract x both side then subtract 2 both side again...what do you get?

OpenStudy (kamille):

ax+2=x-x ax+2-x=0 ax-x=-2

OpenStudy (kamille):

Whats next?:s

OpenStudy (tamtoan):

now factore the x out on the left side and divide the factor to separate x

OpenStudy (kamille):

x(a-2)=-2?

OpenStudy (kamille):

Can you show me?

OpenStudy (tamtoan):

ok divide a-2 to both side of equation

OpenStudy (tamtoan):

what do you get now ?

OpenStudy (kamille):

I don't need to find x, I need to find values of a. But x=-2(a-2)

OpenStudy (tamtoan):

no, if you divide both side by a-2 ...you have x = -2/(a-2) ...and since your solution x is less than 3... x < 3 which is the same as -2/(a-2) < 3 can you solve for a from here?

OpenStudy (kamille):

-2/(a-2)<3 -2-3a+6/a-2<0 -3a+4/(a-2)<0 So a belongs to \[(-\infty;-4/3)U(2;+\infty)\]

OpenStudy (kamille):

Ah, I made a mistake. It was x(a-1). But thanks for your help, anyway : ))

OpenStudy (tamtoan):

i would multiply both side by a-2 (provide a not equal to 2) ...to get -2 < 3(a-2) -2 < 3a - 6 4 < 3a 4/3 < a that would be your a :)

OpenStudy (tamtoan):

your answer have to exclude 2 because a cannot be 2 (denominator can't be zero)